1. The force of repulsion between two positive points charge 5 microfarads and 8microfarads separating at a distance of 0.02m apart is what?
1
Expert's answer
2022-01-18T09:36:09-0500

Given:

"q_1=\rm 5\; \mu C=5*10^{-6}\: C"

"q_2=\rm 8\; \mu C=8*10^{-6}\: C"

"d=0.02\rm\: m"


The Coulomb's law says

"F=k\frac{q_1q_2}{d^2}"

"F=9*10^9*\frac{5*10^{-6}*8*10^{-6}}{(0.02)^2}=900\:\rm N"